1Using Lagrange’s Mean value theorem Prove that log10(x + 1) =

x log10e

(1+θx)

Q.2 Show that excosx = 1 + x + (x2

/2) - (x3

/3) - (11x4

/24)


1
Expert's answer
2021-03-09T02:03:27-0500

1.By mean value theorem:

"f'(c)=\frac{f(b)-f(a)}{b-a},\ c\isin [a,b]"

Let "f(x)=log(x+1)"

Then:

"\frac{1}{x_0+1}=\frac{log(x+1)}{x},\ x_0\isin [0,x],\ x>0"


"log(x+1)=\frac{x}{x_0+1}"

"log(x+1)=xloge^{-(x_0+1)}"

"log10(x+1)=log10+xloge^{-(x_0+1)}"

"log10(x+1)=log10+xloge^{-(\theta x+1)},\ \theta\isin [0,1]"


2.Maclauren series:

"f(x)=\displaystyle\sum_{n=0}^{\infin}\frac{f^{(n)}(0)}{n!}x^n"

Then:

"y(0)=1"

"y'=e^xcosx-e^xsinx=e^x(cosx-sinx)"

"y'(0)=1"

"y''=e^x(cosx-sinx)-e^x(cosx+sinx)=-2e^xsinx"

"y'''=-2e^x(cosx+sinx)"

"y'''(0)=-2"

"y^{(4)}=-2(e^x(cosx+sinx)+e^x(cosx-sinx))=-4e^xcosx"

"y^{(4)}(0)=-4"

"e^xcosx=1+x-\frac{x^3}{3}-\frac{x^4}{6}+..."